Two sides of a triangle have the same length. The third side is 12 cm less than 3 times the length of each equal side. What are the lengths of all three sides if the perimeter is 48 feet?
Answer by mathwiz16(1)

Let x be the equal side and y be the third side.
y = 3x-12
Perimeter = 2x + y = 48 y = 48 - 2x
Make both equations equal:
48 - 2x = 3x - 12
5x = 60
x = 12
Sub in x to solve for y :
y = 3(12) - 12
y = 24